Solve the equation by method of variation of parameters: x^(2)(d^(2)y)/(dx^(2))+x(dy)/(dx)-y=x^(2)e^(x)
1
Expert's answer
2020-11-22T18:46:49-0500

x2y'' +xy' - y=x2ex

homogeneous equation:
x2y'' +xy' - y=0
It can be noted that one of the solutions of the homogeneous equation is

y1=x

Find the second independent solution
according to the Liouville-Ostrogradsky formula:

"W_{y_1y_2}=\begin{vmatrix}
 y_1 & y_2 \\ 
 y_1' & y_2' 
\end{vmatrix} =C_1e^{-\int{\frac{a_1(x)}{a_2(x)}}}"

"y_2'y_1 - y_2y_1'=C_1e^{-\int{\frac{-x}{x^2}dx}}= C_1e^{lnx}=C_1x"

"\frac{y_2'y_1-y_2y_1'}{y_1^2}=\frac{C_1x}{y_1^2}=\frac{C_1x}{x^2}=\frac{C_1}{x}"

"(\frac{y_2}{y_1})'=\frac{C_1}{x}"


"\frac{y_2}{y_1}=C_1ln(x)+C_2"

y2= xC1ln(x)+xC2

Common solution:

y0(x)= xC1ln(x) +xC2

system of equations:

"xC_1'ln(x) + xC_2'=0"

"C_1'(xln(x))' + C_2'(x)'=e^x"


C2'= -C1ln(x)

C1'(ln(x)+1) + C2'= ex


C1'(ln(x) +1) - C1'ln(x) = ex


C1' =ex

C2'= - exln(x)=-1


C1=ex + A1

C2=-x+A2


y(x)= C1(x)xln(x) + C2x= (ex+A1)xln(x) + (-x+A2)x=x+A1xln(x) - x2 +xA2